# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ I.2 Real-World Dataset
Setup. In our real-world dataset experiments, we evaluate performance on two widely used benchmark datasets: TREC Deep Learning (TREC-DL) and NECTAR. The TREC-DL dataset provides 100 candidate answers for each query, offering a rich and diverse set of responses suitable for learning from listwise feedback. In contrast, the NECTAR dataset presents a more concise setup, with only 7 candidate answers per question. From each dataset, we randomly sample $|\mathcal{X}|=5000$ prompts (i.e., questions), each paired with its corresponding set of candidate actions—100 for TREC-DL and 7 for NECTAR.
@@
We use the Gemma-2B language model [78] to construct the feature representation $\phi(x,a)$. To obtain $\phi(x,a)$, we first concatenate the input prompt $x$ and the candidate response $a$ into a single sequence, which is then fed into Gemma-2B. The resulting feature vector is extracted from the last hidden layer of the model and has a dimensionality of $d=2048$. We then apply $\ell_{1}$ normalization to enhance numerical stability and ensure consistent scaling. For each round $t$, we sample the context index from an exponential distribution with rate $\lambda=0.1$, which assigns higher probability to smaller indices and thus biases the selection toward earlier contexts. To generate ranking feedback and evaluate the suboptimality gap, we use the Mistral-7B reward model [33] as the ground-truth reward function, denoted by $r_{\bm{\theta}^{\star}}$.
@@
We measure the suboptimality gap every 2,500 rounds throughout the training process and report the average performance over 10 independent runs, each with a different random seed. Along with the average, we also include the standard error to indicate variability across runs. In these experiments, we report results under the PL loss only, since the performance difference between PL and RB losses is minimal, as demonstrated in the synthetic data experiments. The experiments are conducted on a Xeon(R) Gold 6226R CPU @ 2.90GHz (16 cores) and a single GeForce RTX 3090 GPU.
@@
Baselines. We use the same set of baselines as in the synthetic data experiments. For DopeWolfe [79], we set the sampling size parameter $R$ as $R=\min\{\binom{N}{K},1000\}$. Although a small value of $R\leq 1000$ may introduce significant approximation error—since the theoretically minimal-error choice is $R=\mathcal{O}\big(\binom{N}{K}\big)$—we adopt this smaller value in our experiment to reduce computational overhead.
@@
Figure I.5: Performance comparisons on the TREC-DL dataset (top row) and the NECTAR dataset (bottom row) for varying values of $K=2$, 33, and $5$.
Performance measure. We measure the average realized regret as in the synthetic experiment (Appendix I.1).
@@
Results. We present performance comparisons in Figure I.5. Our algorithm, M-AUPO, consistently outperforms all baselines by a significant margin. As in the synthetic data experiments, the suboptimality gap for all methods decreases as $K$ increases. Notably, DopeWolfe performs particularly poorly on the TREC-DL dataset. This may be attributed to the use of a small sampling size $R$, which is insufficient compared to the full subset space of size $\binom{N}{K}=\mathcal{O}(N^{K})\gg 1000\geq R$. This result highlights an important practical limitation of DopeWolfe: despite its use of approximate optimization to reduce runtime, the method still depends on combinatorial sampling to perform well, which becomes computationally infeasible in large-scale settings. In contrast, our algorithm, M-AUPO, maintains strong performance while requiring only $\tilde{\mathcal{O}}(NK)$ computational cost, making it significantly more scalable and practical for real-world applications.
@@
Table I.3 reports the actual assortment size $|S_{t}|$ selected by M-AUPO on both datasets. In the TREC-DL experiment, $|S_{t}|$ is nearly equal to $K$ for all values of $K$, as the number of available actions is large ($N=100$). In contrast, in the NECTAR experiment, where the number of available actions is much smaller ($N=7$), the actual assortment size $|S_{t}|$ is often smaller than $K$, especially when $K=N$. This reduction occurs because the limited action space constrains the potential informativeness of larger assortments—for example, it becomes difficult to achieve high average uncertainty when there are too few actions to choose from.
@@
TREC-DL dataset, $|S_{t}|$ 2.00 3.00 4.99 6.95
NECTAR dataset, $|S_{t}|$ 2.00 2.99 4.31 4.74
Table I.3: Assortment size $|S_{t}|$ of M-AUPO with varying maximum size $K$ in the real-world dataset experiment

### v2

@@ H.2 Real-World Dataset
Setup. In our real-world dataset experiments, we evaluate performance on two widely used benchmark datasets: TREC Deep Learning (TREC-DL) and NECTAR. The TREC-DL dataset provides 100 candidate answers for each query, offering a rich and diverse set of responses suitable for learning from listwise feedback. In contrast, the NECTAR dataset presents a more concise setup, with only 7 candidate answers per question. From each dataset, we randomly sample $|\mathcal{X}|=5000$ prompts, each paired with its corresponding set of candidate actions—100 for TREC-DL and 7 for NECTAR.
@@
We use the Gemma-2B language model [team2024gemma] to construct the feature representation $\phi(x,a)$. To obtain $\phi(x,a)$, we first concatenate the input prompt $x$ and the candidate response $a$ into a single sequence, which is then fed into Gemma-2B. The resulting feature vector is extracted from the last hidden layer of the model and has a dimensionality of $d=2048$. We then apply $\ell_{1}$ normalization to enhance numerical stability and ensure consistent scaling. For each round $t$, we sample the context index from an exponential distribution with rate $\lambda=0.1$, which assigns higher probability to smaller indices and thus biases the selection toward earlier contexts. To generate ranking feedback and evaluate the suboptimality gap, we use the Mistral-7B reward model [jiang2023mistral7b] as the ground-truth reward function.
@@
We measure the suboptimality gap every 2,500 rounds throughout the training process and report the average performance over 10 independent runs, each with a different random seed. Along with the average, we also include the standard error to indicate variability across runs. In these experiments, we report results under the RB loss only, due to its superior performance, as demonstrated in the synthetic data experiments. The experiments are conducted on a Xeon(R) Gold 6226R CPU @ 2.90GHz (16 cores) and a single GeForce RTX 3090 GPU.
@@
Baselines. We use the same set of baselines as in the synthetic data experiments. For DopeWolfe [thekumparampil2024comparing], we set the sampling size parameter $R$ as $R=\min\{\binom{N}{K},1000\}$. Although a small value of $R\leq 1000$ may introduce significant approximation error—since the theoretically minimal-error choice is $R=\mathcal{O}\big(\binom{N}{K}\big)$—we adopt this smaller value in our experiment to reduce computational overhead.
@@
Figure H.4: Performance comparisons on the TREC-DL dataset (top row) and the NECTAR dataset (bottom row) for varying values of $K=2$, 33, and $5$.
Performance measure. We measure the realized regret as in the synthetic experiment.
@@
Results. We present performance comparisons in Figure H.4. Our algorithm, M-AUPO, consistently outperforms all baselines by a significant margin. As in the synthetic data experiments, the suboptimality gap for all methods decreases as $K$ increases. Notably, DopeWolfe performs particularly poorly on the TREC-DL dataset. This may be attributed to the use of a small sampling size $R$, which is insufficient compared to the full subset space of size $\binom{N}{K}=\mathcal{O}(N^{K})\gg 1000\geq R$. This result highlights an important practical limitation of DopeWolfe: despite its use of approximate optimization to reduce runtime, the method still depends on combinatorial sampling to perform well, which becomes computationally infeasible in large-scale settings. In contrast, our algorithm, M-AUPO, achieves consistently strong performance while incurring only polynomial computational cost, demonstrating superior scalability and practicality for real-world deployment.
@@
Table H.3 reports the actual assortment size $|S_{t}|$ selected by M-AUPO on both datasets. In the TREC-DL experiment, $|S_{t}|$ is nearly equal to $K$ for all values of $K$, as the number of available actions is large ($N=100$). In contrast, in the NECTAR experiment, where the number of available actions is much smaller ($N=7$), the actual assortment size $|S_{t}|$ is often smaller than $K$, especially when $K=N$. This reduction occurs because the limited action space constrains the potential informativeness of larger assortments—for example, it becomes difficult to achieve high average uncertainty when there are too few actions to choose from.
@@
TREC-DL dataset, $|S_{t}|$ 2.00 3.00 5.00 6.98
NECTAR dataset, $|S_{t}|$ 2.00 2.99 4.45 4.76
Table H.3: Assortment size $|S_{t}|$ of M-AUPO with varying $K$ in the real-world dataset experiment